OpenStudy (anonymous):

What is the solution to r over two is equal to one over three? ( r/2=1/3 )

OpenStudy (anonymous):

\[r/2=1/3\] Multiply both sides by 2.\[(2 )(r/2)=(1/3)(2)\] \[r=2/3\]